OpenStudy (anonymous):

Trying to study for a test and need a little help. Please explain as best you can. 16 = -10 + 13x

Nnesha (nnesha):

for this one you have to add 10 both side first

OpenStudy (secret-ninja):

If your purpose is to find x, then your goal should be to get the variable x on one side of the equal side alone. You can start this out by moving the -10 to the other side. Since you are moving it, the sign must change from a -10 to a +10. You are now left with: \[16 + 10 = 13x\]

OpenStudy (secret-ninja):

Now, all you have left is to separate the 13 form the x. Since they are touching, that means they are multiplying, so you must do the opposite (dividing).

OpenStudy (secret-ninja):

You are then left with: \[x = 16 + 10 \div 13\]

OpenStudy (secret-ninja):

Now you just have to solve. 16 + 10 = 26 26 / 13 = 2 So... \[x = 2\]

OpenStudy (secret-ninja):

To check, you just have to insert the 2 for x in the original equation and see if it is still correct. \[16 = -10 + 13 \times 2\] 13 x 2 = 26 26 + (-10) = 16 So it is correct. :)

OpenStudy (secret-ninja):

Did this help? @brittanyluv1796

OpenStudy (secret-ninja):

\(\huge \rlap{\bbox[2.9pt, ,border:6.8px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm \fbox{Welcome}}}}\hspace{127pt}\huge \rlap{\bbox[3.3pt, ,border:7px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm\fbox{to}}}}\) \(\large\color{white}{~~~~~~~~~~~~~~~~~~~~~}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \huge \rlap{\bbox[5pt, ,border:8px solid #2aacd7 ]{\Huge\color{#84c73c}{\rm\fbox{OpenStudy!}}}}\) If this helped, please click the \(\large\sf\color{blue}{Best~Response}\) button and then click the \(\large\sf\color{blue}{Close}\) button to close your question so that you can post a new one later. Go here for all of the basics: http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8 Go Here to learn about the rules of Os: http://openstudy.com/code-of-conduct \(\Huge\cal\color{darkgreen}{Have~a~great~day!}\)

OpenStudy (anonymous):

Yes. Thank you so much. I may have a few more questions. could you help me once i get to them?

OpenStudy (secret-ninja):

Sure, just remember to close this post and open a new one. Once you do, tag me by typing @Secret-Ninja

OpenStudy (anonymous):

please note \[16+10\div 13 \neq \left( 16+10 \right)\div 13\]